How do you solve: 6(x+4)^2=30?

Answer by stanbon(75887)

How do you solve: 6(x+4)^2 = 30
----
Divide both sides by 6 to get:
(x+4)^2 = 24
Take the square root of both sides:
x+4 = +-sqrt(24)
---
x+4 = +-2sqrt(6)
---
x = -4 +-2sqrt(6)
======================
Cheers,
Stan H.